# A quasi-optimal upper bound for induced paths in sparse graphs

### v1

@@ 5 Size of the barriers
Let us denote by $B_{\alpha}(m)$margin: $B_{\alpha}(m)$ the full index-barrier of Definition 3.3 with ${\sf T}:=T_{\alpha}(m)$ as an index-tree. (For simplicity, let us assume that the full index-barrier of the empty tree is defined as the empty word, and that the full index-barrier of $i$ with a single child is defined by $B(i)=i\circ i\circ B(i-1)\circ i\circ i$, i.e., as if $i-1$ was a right child. Note that we can make these assumptions as it upper bounds the size of the actual barrier which is defined as $i\circ B(i-1)\circ i$ in that case.)
@@ 6 The proof of the main theorem
Let $\ell\geq 3$ be some integer. We set ${\sf T}:=T(3,\ell)$, $G:=G({\sf T})$, and $n:=2^{2^{\ell}}$. By Remark 3.17, $G$ has $\Omega(n)$ vertices and by Lemma 3.20 it has a Hamiltonian path, so it has a path of order $\Omega(n)$. By lemma 3.21, $G$ is 222-degenerate. It remains to show that $G$ does not have long induced paths.
@@
$13\cdot\left(\sum_{i=1}^{p}|B(k_{i})|\right)+1\leq 13\cdot\left(|B(\ell)|+\sum_{i=2}^{p}|B(k_{i})|\right)+1$

### v2

@@ 5 Size of the barriers
Let us denote by $B_{\alpha}(m)$margin: $B_{\alpha}(m)$ the full index-barrier of Definition 3.3 with ${\sf T}:=T_{\alpha}(m)$ as an index-tree. (For simplicity, let us assume that the full index-barrier of the empty tree is defined as the empty word, and that the full index-barrier of $i$ with a single child is defined by $B(i)=i\circ i\circ B(i-1)\circ i\circ i$, i.e., as if i1i-1 was a right child. Note that we can make these assumptions as it upper bounds the size of the actual barrier which is defined as $i\circ B(i-1)\circ i$ in that case.)
@@ 6 The proof of the main theorem
Let $\ell\geq 3$ be some integer. We set ${\sf T}:=T(3,\ell)$, $G:=G({\sf T})$, and $n:=2^{2^{\ell}}$. By Remark 3.17, $G$ has $\Omega(n)$ vertices and by Lemma 3.20 it has a Hamiltonian path, so it has a path of order $\Omega(n)$. By lemma 3.21, $G$ is $2$-degenerate. It remains to show that $G$ does not have long induced paths.
@@
$13\cdot\left(\sum_{i=1}^{p}|B(k_{i})|\right)+1\leq 13\cdot\left(|B(\ell)|+\sum_{i=2}^{p}|B(k_{i})|\right)+1.$
@@
Remark 6.1.
We note that the asymptotic dependence in the size of $B(\ell)$ for the order of a longest induced path in our construction is tight, as an induced path may traverse a barrier corresponding to $B(\ell-1,1)$ between a node of rank $\ell-1$ and its parent of rank $1$. Indeed, this barrier has order $\Omega(|B(\ell)|)$. However, it is not clear whether our choice of $\sf T$} results in the smallest barrier, or if the value of $B(\ell)$ is $\Omega(\ell\log\ell)$ for our choice of $\sf T$}.